# R&D Efforts in Cherenkov Imaging Technologies for Particle Identification in Future Experiments

### v1

@@ 3 DIRC detectors at PANDA:
The PANDA experiment [41] at the international accelerator complex FAIR (Facility for Antiproton and Ion Research) in Darmstadt, Germany, will be using high-quality antiproton beam with momenta in the range of 1.5 GeV/c to 15 GeV/c, stored in the HESR (High Energy Storage Ring), to explore fundamental questions of hadron physics in the charmed and multi-strange hadron sector and deliver decisive contributions to the open questions of Quantum ChromoDynamics (QCD). PID plays a central role for the physics program of PANDA. Cherenkov detectors, namely, DIRC detectors in the barrel region and in the end-cap region are cornerstones of the PANDA physics program [61]. Originally inspired by the BaBar DIRC [5] and key results from the R&D activities of the focusing DIRC of the SuperB [57], the PANDA barrel DIRC has brought forward novel ideas by replacing of the water tank expansion volume with compact prisms allowing to operate near magnetic field and insertion of focusing optics for further enhanced PID (see fig. 3 (a) and fig. 3 (b) for main components and schematic of the PANDA barrel DIRC). The compact design of the detector enforces the PANDA DIRC to place photo-sensors in 2 T magnetic field and dense array of anode pixels. The single photon time resolution () should be smaller than 100 ps to allow to compensate the chromatic dispersion effects in the radiator bars and to measure the time-of-propagation of the photons. Moreover, the antiproton–proton average annihilation rate of 20 MHz in the PANDA experiment will lead to densities at the DIRC image planes of 200–800 kHz/cm2 detected photons.
@@
Microchannel Plates (MCPs) are chosen as the baseline photo-sensors for PANDA barrel DIRC thanks to their several advantages; namely, they are commercially available as multi-anode devices and provide a good active area ratio while still being quite compact in size. MCP-PMTs exhibits high gain, high quantum efficiency and good single photon sensitivity, excellent time resolution, much better magnetic field immunity compared to the classical MaPMTs, and low DCR; however they pose severe challenges in terms of aging [1]. In the last decade, MCP-PMTs were reported to show instability in the photocathode performance and loss in quantum efficiency with a Integrated Anode Charge (IAC) of about 200 mC/cm2, whereas the requirement of PANDA remains at a level of 5 C/cm2. A mitigation strategy to lifetime degradation has been a major effort for the PANDA collaboration. It has been demonstrated that improvement in the vacuum quality and sealing enhances the performance. However, an ultra-thin coating of resistive and/or secondary electron emissive (SEE) layer using ALD techniques applied to the surfaces of the MCP pores prevents the desorption of gaseous contaminants during the electron amplification process. An interesting side effect of this approach is the possibility of optimizing the MCP resistance and the SEE coefficient independently and hence increasing the rate capability and the gain of the MCP-PMT. A systematic and thorough characterization of several varieties of MCP-PMTs have demonstrated a ‘tremendously increased lifetime of MCP-PMTs’ (see fig. 3 (c)) with ALD coating [43].
@@
A dedicated beam test exercise has been performed at CERN, with Photonics MCP-PMT with extended lifetime. The results are in excellent agreement with the simulation studies performed with the barrel DIRC. It has been observed that in the polar angle region of 22–140 degrees the prototype provides a separation power of at least 3 for charged particle momenta up to 3.5 GeV/c (see fig. 3 (d)). The final design features narrow radiators made of synthetic fused silica, focusing optics with 3-layer spherical lenses and a compact prism-shaped expansion volume instrumented with MCP-PMTs [32].
@@ 4 Use of PID at the ePIC experiment in EIC:
The ePIC experiment at the EIC is an approved project to be built at Brookhaven National Laboratory (BNL), USA. It aims to address key questions in the field of QCD and hadronic physics. The requirements for the EIC have been described in detail in the common effort by the physics community interested in the EIC physics, known as the Yellow Report [2]. In the Yellow Report, the role of the PID detectors in delivering the physics of the EIC has been considered fundamental. The extensive and challenging PID requirement, like in ePIC, can not be delivered by a single PID technology: in fact, an ample phase-space domain has to be covered including a wide momentum range. Therefore, several technologies have been identified to deliver the PID requirements altogether. The ePIC experiment employs three different Cherenkov-imaging-based technologies: a dual-radiator RICH in the forward direction (1.5 3.5), a proximity-focusing RICH in the backward direction (3.5 1.5), and a high-performance DIRC in the barrel (1.5 1.5) region (fig. 4 (a)). The PID detectors have several challenges [65] and R&D activities related to each of these detectors primarily target solutions to these challenges, risk mitigation and optimization of the detector design [28].
@@ 4.1 The dual-Radiator RICH:
Under these conditions, SiPMs emerge as the only cost-effective solution for covering large active areas, owing to their insensitivity to magnetic fields, high PDE, and excellent timing performance. However, SiPMs exhibit an intrinsic dark count rate (DCR) and are susceptible to radiation damage. Radiation damage leads to an increase in dark current and therefore requires mitigation strategies, such as operating SiPMs at 30C to reduce the DCR and applying periodical high-temperature annealing to recover from radiation damage. In addition, timing information can be exploited using high-precision TDC electronics with gating. Encouraging results on the performance of cooled SiPMs after irradiation and subsequent annealing cycles have been reported [56, 55].
@@
Several beam test studies have been performed with a prototype [60] to validate key features of the dRICH, including radiator optical quality and optical layout. The prototype, in its default configuration, uses Hamamatsu MaPMTs and MAROC chips [20]. The analysis shows that the single-photon resolution for both the aerogel and gas radiators is in excellent agreement with simulation studies. Furthermore, irradiated SiPMs with ALCOR chip–based readout [29] have also been tested with this prototype (see fig. 4 (b) and fig. 4(c) for SiPM arrangements and aerogel ring event displays). A high number of photons were detected, with clear separation between gas and aerogel rings. The measured evolution of the ring radius with particle momentum enabled the identification of electrons, pions, kaons, and protons using aerogel photon information [59]. Initial studies with a pressurized gas radiator have also been carried out to validate the number of detected photons and single-photon resolution. In the following years, a full scale dRICH prototype will be tested at a CERN beam test [51].
@@ 4.2 A high-performance DIRC:
A high-performance DIRC (hpDIRC) technology will provide PID in the ePIC barrel region by distinguishing pions and kaons up to 6 GeV/c with a separation. The hpDIRC is largely inspired by the PANDA DIRC design, as described in Section 3; the ePIC DIRC will reuse the BaBar DIRC bars [5]. The central ideas behind achieving separation at such high momentum are the use of a novel focusing lens system to improve spatial resolution and the exploitation of the timing information of Cherenkov photons. Consequently, the hpDIRC is designed to use commercial MCP-PMTs. Furthermore, refurbishment of the bars, optical gluing of the bars to the expansion volume, and quality control are extremely delicate processes and are altogether progressing promisingly [37].
Dedicated simulation studies have demonstrated that, using realistic parameters of available commercial MCP-PMTs, the hpDIRC can provide the performance required for the ePIC physics program in the presence of a magnetic field. Currently, MCP-PMTs produced by PHOTEK (MAPMT253 series) are assumed as the baseline configuration. However, as a risk-mitigation strategy, DC-coupled HRPPD photo-sensors are also being considered. HRPPD photo-sensors are chosen as the baseline photo-sensors for the pfRICH detector, highlighting the cross-cutting efforts among different subdetector systems within the ePIC community.
The PID performance required for the ePIC hpDIRC is technically challenging, and the targeted upper momentum limit is unprecedented. Nevertheless, through the use of short radiator bars, optimized modular structures, a novel focusing lens system, and fast timing photo-sensors enabling reconstruction in both spatial and temporal domains, the hpDIRC aims to meet its PID requirements. Dedicated beam tests, accompanied by simulation studies and performed in collaboration with the PANDA barrel DIRC group, have shown promising results [37, 38].
@@ 4.3 A Proximity-focusing RICH:
In the electron-going direction, a proximity-focusing RICH (pfRICH) has been chosen as the baseline detector. The ePIC pfRICH is expected to provide separation up to 7 GeV/c with at least 3. Driven by the ePIC physics program, this region is subject to overlapping Cherenkov rings from outgoing jets. Therefore, the pfRICH aims to resolve these rings using not only spatial information but also the timing of Cherenkov photons. The pfRICH departs from the classical proximity-focusing approach by employing a large proximity gap to improve angular resolution ( 45 cm). To increase acceptance, conical mirrors are placed around the vessel wall and beam pipe. This, however, results in complex photon topologies on the sensor surface, particularly for photons emitted near the acceptance boundaries. To address the challenges of overlapping rings and ensure the required performance, exploitation of timing information is central to pfRICH R&D efforts.
Cherenkov photons in the pfRICH will be detected using HRPPDs produced by INCOM, using the same MCP-PMTs employed for LAPPDs [50, 49]. Simulation studies have demonstrated that, with a nominal single–photoelectron time resolution of 50 ps, the pfRICH can achieve separation with high efficiency and low contamination. The photosensors also function as timing detectors. In a dedicated test beam at CERN, gen-II LAPPDs achieved 80 ps time resolution [19]. By exploiting the timing information of Cherenkov photons produced at the detector window by through-going charged particles, the HRPPD sensors can provide precise timing measurements. Thanks to their high quantum efficiency (30%, see fig. 4 (e)), the pfRICH will be capable of detecting an average of 12 photons per track.
The sensors will be placed directly within the acceptance, so the use of SiPMs is ruled out. HRPPDs are subject to loss of gain and PDE in the presence of a magnetic field, particularly at large angles. However, a series of experiments with LAPPDs, a closely related technology produced by the same manufacturer, has demonstrated that both gain and PDE can be recovered by applying a higher bias voltage [7]. Recently, HRPPDs have also been tested in magnetic fields at CERN and BNL. Preliminary results show that HRPPDs present no show-stoppers for pfRICH applications. For readout, a constant-fraction discriminator developed at Fermilab (FCFD) is considered the baseline [68]. Within the EIC community, the AC-LGAD-based TOF in the barrel region will use a similar readout, ensuring a synergistic approach among different subdetectors.
@@ 5 Other relevant R&D activities related to Cherenkov imaging:
The Super Tau Charm Factory (STCF) also requires efficient PID. The barrel PID system is comprised of a proximity focusing RICH, and liquid C6F14 has been considered as a radiator [3]. Hybrid MPGD based photo-sensors based on CsI coated THGEM and Micromegas have been chosen as the default option. The set-up is supposed to provide separation up to 2 GeV/c. Custom built radiator purification, THGEM and Micromegas production, and design of the readout chain are the main directions of their R&D; recently, SiPMs are also being considered as an alternative solution [67].
The MARQ detector at J-PARC hadron facility also aims for a dual radiator RICH with aerogel (n=1.04) and C4F10 gas radiator. The required single photon angular resolution is around 10 mrad. Currently, the hydrophobic aerogel produced by the Aerogel Factory has been chosen as the baseline, and Hamamatsu MPPC S13360-6075CS has been selected as the photo-sensors. To increase the photo-sensitive area, aluminized Mylar foils are considered as light concentrator. Beam test exercises with prototype and ongoing R&D activities are focused on optimizing the light guide, SiPM cooling, and aerogel characteristics [63]. On the other hand, the Belle-II experiment is working on finding an optimal photon detector to replace the HAPD photo-sensors [40]. After 2030 the production of HAPDs are expected to discontinue. MPPC arrays and LAPPD are under consideration. Dedicated R&D efforts related to the sensors and ASIC design is ongoing [42].
@@
| SFC gas alternative | ePIC dRICH, LHCb, MARQ, FCC | ||
| SiPM (MPPC array) | ePIC dRICH, ALICE3, MARQ, possibly FCC, possibly BelleII, CBM upgrade | ||
@@
At CERN, the ALADDIN experiment aims to achieve a precise measurement of the electric dipole moment of charmed baryons ( and ). Suppression of decay products of and is important to ensure efficient reconstruction. To reduce background, ALADDIN plans to use a focusing RICH detector for separating protons from kaons and pions in a range from a few hundred GeV up to the TeV range. This exceptional momentum range requires a large radiator. Currently, ALADDIN RICH is 750 m long. The choice of photo-sensor is critical, simulation studies and R&D activities are ongoing within DRD4 collaboration. Both SiPM and commercial MCP-PMTs are being considered; ALCOR chips are taken as the baseline readout. Depending on the sensor choice, the radiator gas may be He or Ne. He gas may be incompatible with MCP-PMTs due to degassing. The final design choice is extensively studied with R&D efforts, test beams, and dedicated simulation studies [8, 62]. Looking further ahead, the CLD experiment at the FCC-ee aims to use an Array of RICH Cells (ARC), where each cell is a dual radiator RICH. The aerogel radiator serves a radiator and thermal insulator between the cooled SiPM and C4F10 gas. The challenge remains to detect enough photons with an extremely short radiator length ( cm). While the detector is currently validating its proof of principle, it opens several R&D opportunities synergistic to LHCb and ePIC experiments. Dedicated simulation studies are ongoing [25].
6 Conclusion and Acknowledgment

### v2

@@ 3 DIRC detectors at PANDA:
The PANDA experiment [42] at the international accelerator complex FAIR (Facility for Antiproton and Ion Research) in Darmstadt, Germany, will be using a high-quality antiproton beam with momenta in the range of 1.5 GeV/c to 15 GeV/c, stored in the HESR (High Energy Storage Ring), to explore fundamental questions of hadron physics in the charmed and multi-strange hadron sector and deliver decisive contributions to the open questions of Quantum ChromoDynamics (QCD). PID plays a central role in the physics program of PANDA. Cherenkov detectors, namely, DIRC detectors in the barrel region and in the end-cap region, are cornerstones of the PANDA physics program [61]. Originally inspired by the BaBar DIRC [8] and key results from the R&D activities of the focusing DIRC of the SuperB [57], the PANDA barrel DIRC has introduced novel ideas by replacing the water tank expansion volume with compact prisms, allowing it to operate near a magnetic field and insert focusing optics for further enhanced PID (see Fig. 3 (a) for the main components of the PANDA barrel DIRC). The compact design of the detector requires the PANDA DIRC to place photo-sensors in a 2 T magnetic field and with a dense array of anode pixels. The single photon time resolution () should be smaller than 100 ps to compensate for chromatic dispersion effects in the radiator bars and to measure the time-of-propagation of the photons. Moreover, the antiproton–proton average annihilation rate of 20 MHz in the PANDA experiment will lead to densities at the DIRC image planes of 200–800 kHz/cm2 detected photons.
@@
Microchannel Plates (MCPs) are chosen as the baseline photo-sensors for the PANDA barrel DIRC thanks to their several advantages; namely, they are commercially available as multi-anode devices and provide a good active area ratio while still being quite compact in size. MCP-PMTs exhibit high gain, high quantum efficiency, and good single photon sensitivity, excellent time resolution, much better magnetic field immunity compared to the classical MaPMTs, and low DCR; however they pose severe challenges in terms of aging [44]. In the last decade, MCP-PMTs were reported to show instability in the photocathode performance and a loss in quantum efficiency with an Integrated Anode Charge (IAC) of about 200 mC/cm2, whereas the requirement of the PANDA experiment remains at a level of 5 C/cm2. A mitigation strategy to lifetime degradation has been a major effort for the PANDA collaboration. It has been demonstrated that improvements in vacuum quality and sealing enhance the performance. However, an ultra-thin coating of resistive and/or secondary electron emissive (SEE) layer using ALD techniques applied to the surfaces of the MCP pores prevents the desorption of gaseous contaminants during the electron amplification process. An interesting side effect of this approach is the possibility of optimizing the MCP resistance and the SEE coefficient independently and hence increasing the rate capability and the gain of the MCP-PMT. A systematic and thorough characterization of several varieties of MCP-PMTs have demonstrated a ‘tremendously increased lifetime of MCP-PMTs’ with ALD coating [44].
@@
A dedicated beam test exercise has been performed at CERN, with Photonis MCP-PMT with extended lifetime. The results are in excellent agreement with the simulation studies performed with the barrel DIRC. Fig. 3 (c) shows that the log likelihood difference of 7 GeV/c pions from protons is of the order of 5 separation. As direct /K separation was not possible at T9 beam test facility; the observed separation between 7 GeV/c /P was extrapolated to 3.5 GeV/c/K separation, thanks to the similar differences in the Cherenkov angle. The final design features narrow radiators made of synthetic fused silica, focusing optics with 3-layer spherical lenses and a compact prism-shaped expansion volume instrumented with MCP-PMTs [33].
@@ 4 Use of PID at the ePIC experiment in EIC:
The ePIC experiment at the EIC is an approved project to be built at Brookhaven National Laboratory (BNL), USA. It aims to address key questions in the field of QCD and hadronic physics. The requirements for the EIC have been described in detail in the common effort by the physics community interested in the EIC physics, known as the Yellow Report [1]. In the Yellow Report, the role of the PID detectors in delivering the physics of the EIC has been considered fundamental. The extensive and challenging PID requirement, like in ePIC, cannot be delivered by a single PID technology: in fact, an ample phase-space domain has to be covered, including a wide momentum range. Therefore, several technologies have been identified to deliver the PID requirements altogether. The ePIC experiment employs three different Cherenkov-imaging-based technologies: a dual-radiator RICH in the forward direction (1.5 3.5), a proximity-focusing RICH in the backward direction (3.5 1.5), and a high-performance DIRC in the barrel (1.5 1.5) region (Fig. 4 (a)). The PID detectors have several challenges and R&D activities related to each of these detectors primarily target solutions to these challenges, risk mitigation and optimization of the detector design [65].
@@ 4.1 The dual-radiator RICH:
Under these conditions, SiPMs emerge as the only cost-effective solution for covering large active areas, owing to their insensitivity to magnetic fields, high PDE, and excellent timing performance. However, SiPMs exhibit an intrinsic DCR and are susceptible to radiation damage. Radiation damage leads to an increase in dark current and therefore requires mitigation strategies, such as operating SiPMs at 30C to reduce the DCR and applying periodical high-temperature annealing to recover from radiation damage. In addition, timing information can be exploited using high-precision TDC electronics with gating. Encouraging results on the performance of cooled SiPMs after irradiation and subsequent annealing cycles have been reported [56, 55].
@@
Several beam test studies have been performed with a prototype [66] to validate key features of the dRICH, including radiator optical quality and optical layout. The prototype, in its default configuration, uses Hamamatsu MaPMTs and MAROC chips [22]. The analysis shows that the single-photon resolution for both the aerogel and gas radiators is in excellent agreement with simulation studies. Furthermore, irradiated SiPMs with ALCOR chip–based readout [30] have also been tested with this prototype (see Fig. 4 (c) for SiPM arrangements). A high number of photons were detected, with clear separation between gas and aerogel rings. The measured evolution of the ring radius with particle momentum enabled the identification of electrons, pions, kaons, and protons using aerogel photon information [59]. Initial studies with a pressurized gas radiator have also been carried out to validate the number of detected photons and single-photon resolution. In the following years, a full scale dRICH prototype will be tested at CERN [51].
@@ 4.2 A high-performance DIRC:
A high-performance DIRC (hpDIRC) technology will provide PID in the ePIC barrel region by distinguishing pions and kaons up to 6 GeV/c with a separation. The hpDIRC is largely inspired by the PANDA DIRC design, as described in Section 3; the ePIC DIRC will reuse the BaBar DIRC bars [8]. The central ideas behind achieving separation at such high momentum are the use of a novel focusing lens system to improve spatial resolution and the exploitation of the timing information of Cherenkov photons. Consequently, the hpDIRC is designed to use commercial MCP-PMTs. Furthermore, refurbishment of the bars, optical gluing of the bars to the expansion volume, and quality control are extremely delicate processes and are altogether progressing promisingly.
Dedicated simulation studies have demonstrated that, using realistic parameters of available commercial MCP-PMTs, the hpDIRC can provide the performance required for the ePIC physics program in the presence of a magnetic field. Currently, MCP-PMTs produced by PHOTEK (MAPMT253 series) are assumed as the baseline configuration. However, as a risk-mitigation strategy, DC-coupled HRPPD [49] photo-sensors are also being considered. HRPPD photo-sensors are chosen as the baseline photo-sensors for the pfRICH detector, highlighting the cross-cutting efforts among different subdetector systems within the ePIC community.
The PID performance required for the ePIC hpDIRC is technically challenging, and the targeted upper momentum limit is unprecedented for a detector of this type. Nevertheless, through the use of short radiator bars, optimized modular structures, a novel focusing lens system, and fast timing photo-sensors enabling reconstruction in both spatial and temporal domains, the hpDIRC aims to meet its PID requirements. Dedicated beam tests, accompanied by simulation studies and performed in collaboration with the PANDA barrel DIRC group, have shown promising results [39].
@@ 4.3 A Proximity-focusing RICH:
In the electron-going direction, a proximity-focusing RICH (pfRICH) has been chosen as the baseline detector. The ePIC pfRICH is expected to provide at least 3 separation up to 7 GeV/c. Driven by the ePIC physics program, this region is subject to overlapping Cherenkov rings from outgoing jets. Therefore, the pfRICH aims to resolve these rings using not only spatial information but also the timing of Cherenkov photons. The pfRICH departs from the classical proximity-focusing approach by employing a large proximity gap to improve angular resolution ( 45 cm). To increase acceptance, conical mirrors are placed around the vessel wall and beam pipe. This, however, results in complex photon topologies on the sensor surface, particularly for photons emitted near the acceptance boundaries. To address the challenges of overlapping rings and ensure the required performance, the exploitation of timing information is central to pfRICH R&D efforts.
Cherenkov photons in the pfRICH will be detected using HRPPDs produced by INCOM, using the same MCP-PMTs employed for LAPPDs [50, 49]. Simulation studies have demonstrated that, with a nominal single–photoelectron time resolution of 50 ps, the pfRICH can achieve separation with high efficiency and low contamination. The photosensors also function as timing detectors. In a dedicated test beam at CERN, gen-II LAPPDs achieved 80 ps time resolution [21]. By exploiting the timing information of Cherenkov photons produced at the detector window by through-going charged particles, the HRPPD sensors can provide precise timing measurements. Thanks to their high quantum efficiency (30% [49]), the pfRICH will be capable of detecting an average of 12 photons per track.
The HRPPDs are subject to loss of gain and PDE in the presence of a magnetic field, particularly at large angles. However, a series of experiments with LAPPDs, a closely related technology produced by the same manufacturer, have demonstrated that both gain and PDE can be recovered by applying a higher bias voltage [10]. Recently, HRPPDs have also been tested in magnetic fields at CERN and BNL. Preliminary results show that HRPPDs present no show-stoppers for pfRICH applications. For readout, a constant-fraction discriminator developed at Fermilab (FCFD) is considered as the baseline [69]. Within the EIC community, the AC-LGAD-based TOF in the barrel region will use a similar readout, ensuring a synergistic approach among different subdetectors.
@@ 5 Other relevant R&D activities related to Cherenkov imaging:
The Super Tau Charm Factory (STCF) also requires efficient PID. The barrel PID system is comprised of a proximity focusing RICH, and liquid C6F14 has been considered as a radiator [6]. Hybrid Micro Pattern Gaseous Detector (MPGD) [60] based photo-sensors based on CsI coated THick Gas Electron Multiplier (THGEM) [23] and Micromegas [34] have been chosen as the default option. The set-up is supposed to provide separation up to 2 GeV/c. Custom built radiator purification, THGEM and Micromegas production, and design of the readout chain are the main directions of their R&D; recently, SiPMs are also being considered as an alternative solution [68].
The J-PARC hadron facility also aims for a dual radiator RICH with aerogel (n=1.04) and C4F10 gas radiator [63]. The required single photon angular resolution is around 10 mrad. Currently, the hydrophobic aerogel produced by the Aerogel Factory has been chosen as the baseline, and Hamamatsu MPPC S13360-6075CS has been selected as the photo-sensors. To increase the photo-sensitive area, aluminized Mylar foils are considered as light concentrators. Beam test exercises with prototype and ongoing R&D activities are focused on optimizing the light guide, SiPM cooling, and aerogel characteristics [63]. On the other hand, the Belle-II experiment is working on finding an optimal photon detector to replace the HAPD photo-sensors [41]. After 2030 the production of HAPDs are expected to discontinue. Multi-Pixel Photon Counter (MPPC) arrays and LAPPDs are under consideration. Dedicated R&D efforts related to the sensors and ASIC design are ongoing [43].
@@
| SFC gas alternative | ePIC dRICH, LHCb, J-PARC, FCC | ||
| SiPM (MPPC array) | ePIC dRICH, ALICE3, LHCb, J-PARC, possibly FCC, possibly BelleII, CBM upgrade | ||
@@
At CERN, the ALADDIN experiment [11] aims to achieve a precise measurement of the electric dipole moment of charmed baryons ( and ). Suppression of decay products of and is important to ensure efficient reconstruction. To reduce background, ALADDIN plans to use a focusing RICH detector for separating protons from kaons and pions in a range from a few hundred GeV up to the TeV range. This exceptional momentum range requires a large radiator. Currently, the ALADDIN RICH is 750 cm long. The choice of photo-sensor is critical, simulation studies and R&D activities are ongoing within the DRD4 collaboration. Both SiPM and commercial MCP-PMTs are being considered; ALCOR chips are taken as the baseline readout. Depending on the sensor choice, the radiator gas may be He or Ne. He gas may be incompatible with MCP-PMTs due to degassing. The final design choice is extensively studied with R&D efforts, test beams, and dedicated simulation studies [62]. Looking further ahead, the CLD experiment at the FCC-ee aims to use an Array of RICH Cells (ARC), where each cell is a dual radiator RICH [27]. The aerogel radiator serves both as a radiator and a thermal insulator between the cooled SiPM and C4F10 gas. The challenge remains to detect enough photons with an extremely short radiator length ( cm). While the detector is currently validating its proof of principle, it opens several R&D opportunities synergistic to the LHCb and ePIC experiments. Dedicated simulation studies are ongoing.
Recent tests with the prototype RICH of the CBM experiment demonstrated the compatibility of the DiRICH with the novel DAQ concept developed. In the baseline design, the MaPMT (H12700) is used. A future upgrade with SiPM photo-sensors has been considered and subject to ongoing R&D activities [54, 19].
6 Conclusion and Acknowledgments